OpenStudy (anonymous):

Find the length of the major axis of the ellipse with equation 9(x+4)^2+4(y-1)^2=36 A)4 B)18 C)6 D)9

OpenStudy (amistre64):

the 9 and 4 give it away since we got a 36 over there; so the unders go to 9 and 4; since 9 is bigger sqrt it and double it

OpenStudy (anonymous):

so the answer is 9?

OpenStudy (amistre64):

no, the answer is based on the 9

OpenStudy (amistre64):

2sqrt(9)

OpenStudy (anonymous):

so its 4

OpenStudy (amistre64):

maybe ....

OpenStudy (anonymous):

i think im gonna go with 4

OpenStudy (anonymous):

it may be 3

OpenStudy (amistre64):

i wouldnt go with 4 ...

OpenStudy (anonymous):

3 isnt an option lol

OpenStudy (amistre64):

3 is closer than 4 ;)

OpenStudy (anonymous):

i dont understand lol

OpenStudy (amistre64):

2 sqrt(9) doesnt equal 4

OpenStudy (anonymous):

18

OpenStudy (amistre64):

you need new batteries in your calculator i think

OpenStudy (anonymous):

18 isnt right either

OpenStudy (amistre64):

nope, it aint so, it aint 9, it aint 4, and it aint 18

OpenStudy (anonymous):

\[\left( \frac{x+3}{2} \right)^2+\left( \frac{y-1}{3} \right)^2=1\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\left( \frac{x+4}{2} \right)^2+\left( \frac{y-1}{3} \right)^2=1\]

OpenStudy (anonymous):

it should be 3

OpenStudy (anonymous):

sorry 2*3=6 major